# ProLLaMA: A Protein Large Language Model for Multi-Task Protein Language Processing

### v1

@@ Abstract
Large Language Models (LLMs), including GPT-x and LLaMA2, have achieved remarkable performance in multiple Natural Language Processing (NLP) tasks. Under the premise that protein sequences constitute the protein language, Protein Large Language Models (ProLLMs) trained on protein corpora excel at de novo protein sequence generation. However, as of now, unlike LLMs in NLP, no ProLLM is capable of multiple tasks in the Protein Language Processing (PLP) field. This prompts us to delineate the inherent limitations in current ProLLMs: (i) the lack of natural language capabilities, (ii) insufficient instruction understanding, and (iii) high training resource demands. To address these challenges, we introduce a training framework to transform any general LLM into a ProLLM capable of handling multiple PLP tasks. Specifically, our framework utilizes low-rank adaptation and employs a two-stage training approach, and it is distinguished by its universality, low overhead, and scalability. Through training under this framework, we propose the ProLLaMA model, the first known ProLLM to handle multiple PLP tasks simultaneously. Experiments show that ProLLaMA achieves state-of-the-art results in the unconditional protein sequence generation task. In the controllable protein sequence generation task, ProLLaMA can design novel proteins with desired functionalities. In the protein property prediction task, ProLLaMA achieves nearly 100% accuracy across many categories. The latter two tasks are beyond the reach of other ProLLMs. Code is available at https://github.com/Lyu6PosHao/ProLLaMA.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs), like GPT-x and LLaMA2 (Bubeck et al., 2023; Touvron et al., 2023), have achieved outstanding performance in handling a wide range of Natural Language Processing (NLP) tasks (Tamkin et al., 2021; Zhao et al., 2023; Kocoń et al., 2023; Qin et al., 2023; Huang et al., 2023; Zhong et al., 2023; Bang et al., 2023), including both Natural Language Generation (NLG) and Natural Language Understanding (NLU) tasks. This surge in LLMs has extended their applications beyond traditional contexts, including their adoption in the challenging field of protein design (Notin et al., 2022; Strokach & Kim, 2022; Ferruz & Höcker, 2022).
Figure 1: Analogous to general LLMs managing multiple tasks in the NLP field, our ProLLaMA can handle multiple tasks in Protein Language Processing. In contrast, other ProLLMs are limited to the single task of unconditional protein generation.
@@
Taking protein sequences as the protein language, researchers train models with architectures similar to current LLMs on a vast protein corpus (Strokach & Kim, 2022; Ferruz et al., 2022; Nijkamp et al., 2023; Moffat et al., 2022; Madani et al., 2023; Qin et al., 2023). These Protein LLMs (ProLLMs) have paved the way for the rapid generation of structurally plausible protein sequences, holding immense potential for biomedical and biotechnological innovations (Notin et al., 2022; Nijkamp et al., 2023). However, this progress is met with challenges, particularly in extending their capabilities beyond sequence generation.
@@
Analogous to NLP, tasks related to protein language can be viewed as Protein Language Processing (PLP) (Bepler & Berger, 2021; Ofer et al., 2021). It is evident that current ProLLMs primarily focus on a single task in PLP, not covering multiple tasks like LLMs do in NLP (Ferruz et al., 2022; Madani et al., 2023). These limitations prompt the need for innovative solutions to unleash the full potential of ProLLMs. Developing a multi-tasking ProLLM would be highly beneficial for protein engineering and understanding the protein fitness landscape (Wright et al., 1932; Pan & Kortemme, 2021; Ren et al., 2022; Song & Li, 2023), but three main challenges must be considered:
@@
(i) Necessity of Natural Language: Protein language is not fully sufficient for PLP tasks, meaning it cannot fully represent all components of a task (the user instruction and the expected output) (Xu et al., 2023; Wang et al., 2023). The instruction and the output require a language beyond protein language (typically, natural language) for representation, which current ProLLMs lack.
@@
(ii) Instruction Following: To possess multi-tasking capabilities, models must execute tasks following user instructions (Zeng et al., 2023; Liu et al., 2023b; Zhou et al., 2023). However, current ProLLMs are unable to follow instructions.
@@
(iii) Training Resource Consumption: Substantial training resources are needed for models to learn natural language, protein language, and user instructions (Cui et al., 2023), which can sometimes be unaffordable.
@@
To address the challenges, we propose a two-stage training framework to achieve a ProLLM for multi-task PLP. In the first stage, we leverage a pre-trained general LLM like LLaMA2 to continually learn the protein language while maintaining the natural language knowledge of the model. In the second stage, the model is further trained on a multi-task PLP dataset through instruction tuning to equip the model with instruction following capabilities and multi-task capabilities of PLP. Notably, during both stages, we adopt Low-Rank Adaptation (LoRA) (Hu et al., 2021), which prevents catastrophic forgetting in the first stage, enhances scalability in the second stage and reduces the training cost.
@@
Using the training framework, we develop ProLLaMA, a model capable of multi-tasks for PLP, distinguishing it from all other ProLLMs. Through a series of experiments, we demonstrate the multitasking capabilities of our ProLLaMA. Specifically, in unconditional protein generation, ProLLaMA outperforms current ProLLMs and other sequence-generative models on common metrics such as pLDDT and TM-score. In controllable protein generation, based on a user-provided textual description of the desired protein functionalities, ProLLaMA generates novel proteins from scratch with functionalities as competent as natural proteins, such as SAM-MT and Trx. For protein property prediction, ProLLaMA achieves an average accuracy of 72% on the test dataset and nearly 100% accuracy in many subcategories. In summary, the contributions of our research are as follows:
@@
$\bullet$ We propose a training framework with scalability and efficiency that enables any general LLM to be trained as a proficient model for multiple tasks in Protein Language Processing.
$\bullet$ Our ProLLaMA is the first model to our knowledge capable of simultaneously handling multiple PLP tasks.
@@
$\bullet$ Experiments show that our ProLLaMA not only handles PLP tasks beyond the reach of existing ProLLMs but also achieves state-of-the-art results in the protein generation task where current ProLLMs are active.
@@ 2 Preliminaries
2.1 Necessity of Natural Language
@@
As aforementioned, given the similarities between protein sequences and natural language, tasks related to protein sequences can be considered as Protein Language Processing (PLP), analogous to NLP. However, we observe a fundamental difference between protein language and natural language: natural language is complete for NLP tasks, whereas protein language is not complete for PLP tasks.
Figure 2: The overall architecture of our ProLLaMA. We add low-rank adapters (LoRA) to certain weights. During the training process, we freeze these weights and other parameters, focusing solely on training LoRA.
Specifically, natural language can represent all components of a task (i.e., the input instructions and the expected output) (Liu et al., 2023a). For example, in a sentiment analysis task, the instruction and output can be expressed as “Analyze the sentiment contained in this sentence: I am happy” and “The sentiment is positive” respectively, all in natural language. However, for PLP tasks, the instructions and outputs cannot be fully represented in protein language. For instance, in the protein property prediction task, the task instruction could be ”Predict the property of this protein: MAFCF…FEV”, with the expected output being ”The property is Trx superfamily.” It is evident that both the instructions and outputs of tasks require assistance from a language beyond protein language, in this case, natural language, for representation.
Therefore, multi-task ProLLMs must possess a certain level of natural language ability, especially as more textual descriptions of proteins become available (Xu et al., 2023).
2.2 Causal Language Modeling
Causal Language Modeling (CLM) is the objective of training LLMs. Given a token sequence $\boldsymbol{x}=\{x_{0},x_{1},\ldots,x_{n-1}\}$ that is subject to training. CLM can be conceptualized as predicting the $i$-th token based on the preceding $i-1$ tokens (Bengio et al., 2000). Therefore, the optimization objective when training LLM using CLM is formulated as:
$\small\mathcal{L}(\Theta)=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D}}\left[-\sum_{i}\log p(x_{i}|x_{0},x_{1},\ldots,x_{i-1};\Theta)\right]$ (1)
2.3 Low Rank Adaptation
@@
Low-Rank Adaptation (LoRA) (Hu et al., 2021) is a parameter-efficient technique for fine-tuning LLMs. Due to the immense parameter size of LLMs, full-parameter fine-tuning could be impractical sometimes. LoRA circumvents this by freezing the original parameters of LLMs and introducing additional trainable low-rank adapters. It achieves fine-tuning with a significantly smaller number of trainable parameters, yielding results comparable to full-parameter fine-tuning.

### v2

# ProLLaMA: A Protein Language Model for Multi-Task Protein Language Processing

@@ Abstract
Large Language Models (LLMs) have achieved remarkable performance in multiple Natural Language Processing (NLP) tasks. Under the premise that protein sequences constitute the protein language, Protein Language Models(PLMs) have advanced the field of protein engineering. However, as of now, unlike LLMs in NLP, PLMs cannot handle the protein understanding task and the protein generation task simultaneously in the Protein Language Processing (PLP) field. This prompts us to delineate the inherent limitations in current PLMs: (i) the lack of natural language capabilities, (ii) insufficient instruction understanding, and (iii) high training resource demands. To address these challenges, we introduce a training framework to transform any general LLM into a PLM capable of handling multiple PLP tasks. To improve training efficiency, we propose Protein Vocabulary Pruning (PVP) for general LLMs. We construct a multi-task instruction dataset containing 13 million samples with superfamily information, facilitating better modeling of protein sequence-function landscapes. Through these methods, we develop the ProLLaMA model, the first known PLM to handle multiple PLP tasks simultaneously. Experiments show that ProLLaMA achieves state-of-the-art results in the unconditional protein sequence generation task. In the controllable protein sequence generation task, ProLLaMA can design novel proteins with desired functionalities. As for the protein understanding task, ProLLaMA achieves a 62% exact match rate in superfamily prediction. Codes, model weights, and datasets are available at https://github.com/PKU-YuanGroup/ProLLaMA and https://huggingface.co/GreatCaptainNemo.
@@ 1 Introduction
Large Language Models (LLMs), like GPT-x and LLaMA2 [1, 2], have achieved outstanding performance in handling a wide range of Natural Language Processing (NLP) tasks [3, 4, 5, 6, 7, 8, 9], including both Natural Language Generation (NLG) and Natural Language Understanding (NLU) tasks, in a generative manner. This surge in LLMs has extended their applications beyond traditional contexts, including their adoption in the challenging field of protein engineering [10, 11, 12, 13, 14].
@@
Taking protein sequences as the protein language, researchers train Protein Language Models (PLMs) on vast protein corpora [6, 11, 12]. PLMs have the potential to significantly advance protein engineering, holding immense promise for biomedical and biotechnological innovations [10]. However, this progress is challenged, particularly in extending their capabilities to multi-task Protein Language Processing (PLP).
Analogous to NLP, tasks related to protein language can be viewed as PLP [15, 16]. Consequently, PLP tasks can also be divided into two categories: Protein Language Generation (PLG), such as protein sequence generation, and Protein Language Understanding (PLU), such as protein property prediction. However, current PLMs tend to focus either exclusively on PLG [17, 18, 19] or PLU [20, 21, 22, 23, 24], rather than being proficient in both aspects simultaneously, as LLMs in NLP.
@@
These limitations prompt the need for innovative solutions to unleash the full potential of PLMs. Developing a multi-task PLM would be highly beneficial for protein engineering and protein fitness landscape modeling [25, 26, 27, 28], but three main challenges must be considered:
@@
(i) Necessity of Natural Language: Protein language is not fully sufficient for PLP tasks, meaning it cannot fully represent all components of a task (the task instruction, the input, and the expected output) [29, 30]. It requires a language beyond protein language (typically, natural language) for representation, which current PLMs lack.
@@
(ii) Instruction Following: To possess multi-task capabilities, models must execute tasks following user instructions [31, 32, 33]. However, current PLMs are unable to follow instructions.
@@
(iii) Training Resource Consumption: Substantial training resources are needed for models to learn natural language, protein language, and user instructions [34], which can sometimes be unaffordable.
@@
To address the challenges, we construct an instruction dataset that contains approximately 13 million samples and encompasses both PLG and PLU tasks. We propose a two-stage training framework to achieve a PLM for multi-task PLP. In the first stage, we leverage a pre-trained general LLM like LLaMA2 to continually learn the protein language while maintaining the natural language knowledge. In the second stage, the model is further trained on the multi-task instruction dataset. Additionally, we propose Protein Vocabulary Pruning (PVP) for general LLMs to significantly improve training efficiency. Furthermore, during both stages, we adopt Low-Rank Adaptation (LoRA) [35], which prevents catastrophic forgetting in the first stage and reduces training costs in both stages.
Figure 1: Left: LLMs can handle both generation and understanding tasks, whereas PLMs cannot. This highlights the disparity in capabilities between the two. Right: Our ProLLaMA can handle generation tasks (unconditional protein generation, controllable protein generation) and understanding tasks (protein superfamily prediction), surpassing current PLMs.
@@
Using these methods, we develop ProLLaMA, a model capable of multi-tasks for PLP, distinguishing it from all other PLMs. Through a series of experiments, we demonstrate the multi-task capabilities of our ProLLaMA. Specifically, as for PLG, in unconditional protein generation, ProLLaMA outperforms current PLMs on common metrics such as pLDDT and TM-score. In controllable protein generation, based on a user-provided textual description, ProLLaMA generates novel proteins from scratch with desired functionalities, such as the SAM-MT superfamily. As for PLU, in protein superfamily prediction, ProLLaMA achieved a 62% exact match rate on the test dataset and obtained an F1-score above 0.9 in many specific categories. In summary, the contributions of our research are as follows:
@@
$\bullet$ We propose a training framework that enables any general LLM to be trained as a proficient model for multi-task PLP, including both PLG and PLU tasks.
$\bullet$ We propose Protein Vocabulary Pruning and are the first to apply Low-Rank Adaptation in the training of large PLMs, which improves the efficiency of protein learning.
$\bullet$ We construct an instruction dataset that contains 13 million samples and over 11,000 kinds of superfamily annotations, potentially facilitating a better modeling of sequence-function landscapes.
@@
$\bullet$ Experiments show that our ProLLaMA not only handles multiple PLP tasks but also achieves state-of-the-art results in protein generation tasks.
@@ 2 Preliminaries
Figure 2: (A) Overview of the dataset construction. The protein language dataset contains 53 million samples, which is used for training in Stage 1. The instruction dataset contains 13 million instances with 11,268 unique superfamily annotations, which is used for training in Stage 2. (B) Overview of the training framework. Stage 1: The pre-trained LLaMA2 learns the protein language, resulting in ProLLaMA. Stage 2: ProLLaMA learns to perform multiple tasks by instruction tuning. (C) Overview of protein vocabulary pruning. The vocabulary, Embed layer and Head layer are pruned based on characteristics of protein datasets, aiming for higher training efficiency.
@@
Necessity of Natural Language. As aforementioned, given the similarities between protein sequences and natural language, tasks related to protein sequences can be considered as PLP, analogous to NLP. However, we observe a fundamental difference between protein language and natural language: natural language is complete for NLP tasks, whereas protein language is not complete for PLP tasks.
For example, in the sentiment analysis task, instructions, inputs, and outputs are straightforwardly expressed in natural language, such as “Analyze the sentiment: I am happy” yielding “The sentiment is positive.” However, in PLP tasks like protein property prediction, instructions like “Predict this protein’s property: MAFCF…FEV” cannot be fully conveyed in protein language alone, necessitating assistance from a language beyond protein language, in this case, natural language, for representation. Therefore, multi-task PLMs must possess a certain level of natural language ability, especially as more textual descriptions of proteins become available [29].
@@
Low-Rank Adaptation (LoRA) [35] is a parameter-efficient technique for fine-tuning LLMs. Due to the immense parameter size of LLMs, full-parameter fine-tuning could be impractical sometimes. LoRA circumvents this by freezing the original parameters of LLMs and introducing additional trainable low-rank adapters. It achieves fine-tuning with a significantly smaller number of trainable parameters, yielding results comparable to full-parameter fine-tuning. LoRA prevents catastrophic forgetting of the original knowledge, as the newly learned knowledge has a lower rank than the original knowledge. The theoretical details of LoRA are provided in Appendix A.2.
3 Methods
Figure 3: The overview of the ProLLaMA model. We add low-rank adapters (LoRA) to certain weights and apply protein vocabulary pruning (PVP). We freeze original parameters, focusing solely on training LoRA (Embed and Head are also involved in the first training stage).
In Section 3.1 and Figure 2(A), we show how to construct the protein language dataset and the instruction dataset. In Section 3.2 and Figure 2(C), we show how Protein Vocabulary Pruning (PVP) works. In Section 3.3, we show how the LLaMA2 model learns protein language on the protein language dataset through continued pre-training, resulting in ProLLaMA. In Section 3.4, we show the integration of various tasks into ProLLaMA through instruction tuning on the instruction dataset. The overview of the training framework is shown in Figure 2(B). The overview of the ProLLaMA model is shown in Figure 3.
3.1 Dataset Construction
The protein language dataset is utilized in the first training stage to enable LLaMA2 to grasp the language of proteins. Specifically, the dataset is sourced from UniRef50_2023_03 [36] on the UniProt website. We eliminate the descriptive parts of UniRef50, retaining only the pure protein sequences. Furthermore, We filter UniRef50 to ensure that the protein sequences consisted only of the 20 standard amino acids. We also retain sequences with a length of less than 512, aligning with ProGen [18]. Given that the lengths of protein sequences follow a long-tail distribution, the sequences that are deleted constitute only a small portion of the total dataset. To preprocess the protein sequences, we employ a specific prefix “Seq=<” and suffix “>”. This standardized format aids LLaMA2 in distinguishing the new protein language from its existing natural language knowledge, thereby reducing confusion. The original Uniref50 contains 60,952,894 sequences, while after the above series of processing, our dataset comprises 52,807,283 protein sequences, with 90% for training and 10% reserved for testing.
The instruction dataset is utilized in the second training stage to enable ProLLaMA to perform various tasks. We first obtain the protein2ipr database from InterPro [37], which includes all proteins from UniProtKB along with their corresponding InterPro annotation information. Subsequently, we iterate through each protein’s rep_id in UniRef50 to retrieve the corresponding annotation information from protein2ipr. This retrieval process is implemented using a distributed Redis database, and only proteins with lengths less than 256 participate in the retrieval to enhance efficiency. We utilize regular expressions to filter out superfamily annotation from the whole annotation. In the end, we obtain 6,350,106 data instances, each of which contains one protein sequence and its superfamily annotation. And the number of unique superfamily annotations is 11,268.
Then, we process the obtained data into a multi-task instruction dataset following the Alpaca format [38], where each instance comprises three parts: instruction, input, and output. The instruction specifies the task type. We design two tasks: generating proteins based on superfamily and determining the superfamily of the given protein. For the former task, the input is the superfamily annotation, and the output is the expected protein. The latter task is the opposite. We selected these two tasks because they represent PLG and PLU, respectively. Additionally, modeling the relationships between sequences and superfamilies is crucial for uncovering protein functions and evolutionary insights [39]. In the end, the instruction dataset comprises 12,700,212 (6,350,106 * 2) instances, with 90% for training and the rest reserved for testing.
3.2 Protein Vocabulary Pruning
It is known that LLMs are equipped with large vocabularies in tokenizers, which is beneficial to cover multi-lingual corpora and shorten the length of the input token sequence. However, a larger vocabulary also implies slower tokenization of raw text, along with increased parameters in the model’s Embed layer and Head layer, as discussed in detail in Appendix A.9. When considering only the protein language, such large vocabularies are clearly unnecessary. Therefore, we propose a method called Protein Vocabulary Pruning (PVP). PVP helps ProLLaMA achieve almost the same performance with higher training efficiency, whose proof is in Appendix A.3.
Table 1: Effects of PVP. The vocabulary size, parameter numbers in Embed and Head layers, parameter numbers involved in training are listed.
Training Stage Vocab Embed and Head Training
I (w/o PVP) 32,000 262M 582M
I (w/ PVP) 1,045 8.56M 328M
II (w/o PVP) 32,000 262M 160M
II (w/ PVP) 25,466 209M 160M
PVP is rule-based. Specifically, we iterate through each token in the original vocabulary and retain it if it meets our predefined grammatical rules (e.g., consisting only of the 20 uppercase English letters representing standard amino acids). We save these tokens and their indices. Using the saved tokens, we construct a smaller vocabulary. We then use the indices to extract the corresponding vectors from Embed and Head layers to construct smaller ones. Thus, we can train LLMs using reduced vocabulary and parameters.
PVP is employed before training. After training, the vocabulary can be recovered back to its original size by incorporating the entries from the original vocabulary. This process is also applicable to the Embed and Head layers, restoring them to their original dimensions. This ensures that the architecture of LLMs remains entirely consistent both before and after training, facilitating potential training on broader corpora in the future. Naturally, if subsequent training is not a consideration, omitting the recovery operation is permissible.
We use PVP in both the first and second training stages of ProLLaMA, and its effects are shown in Table 1. It is evident that we achieve a significant compression rate in the first stage, primarily because the dataset in the first stage contains only protein sequences, which can be summarized by a simple rule. In contrast, the dataset in the second stage also includes English. More details about PVP is shown in Appendix A.3.
3.3 Learning Protein Language
As mentioned in Section 1, current PLMs lack natural language abilities, which hinders multi-task capabilities. To solve this problem, we propose leveraging LLaMA2 to perform continued pre-training on protein language. This approach is analogous to humans learning a foreign language, where the model learns protein language while retaining its original natural language abilities.
We add Low-Rank Adapters (LoRA) into LLaMA2. To be specific, in each decoder block, we add LoRA to certain weights including $\textit{W}_{q}$, $\textit{W}_{k}$, $\textit{W}_{v}$, $\textit{W}_{o}$, $\textit{W}_{up}$, $\textit{W}_{gate}$ and $\textit{W}_{down}$. The original parameters are frozen, enabling only LoRA to be trained. Due to the significant differences between protein language and natural language, we choose a relatively high rank for LoRA, which helps the model learn protein sequences better and prevents under-fitting. We also include both the Embed and Head layers in training. This is based on the premise that a token may have different meanings in protein sequences and natural languages, requiring distinct embeddings for the same token.
We trained the model using causal language modeling on the protein language dataset, resulting in ProLLaMA. Details on causal language modeling are provided in Appendix A.1. Benefiting from PVP and LoRA, as shown in Table 7, we train only about 5% of the parameters, in contrast to full-parameter training, which significantly reduces training costs. Additionally, as the remaining parameters are not involved in training, the inherent natural language abilities are preserved.
In summary, we have developed ProLLaMA, a model that comprehends both protein language and natural language, with reduced training costs. Consequently, we have addressed two problems mentioned in Section 1: the lack of natural language abilities and excessive training costs.
3.4 Performing Multiple Tasks
As aforementioned, current PLMs are unable to perform multiple tasks based on user instructions. To solve this problem, we perform instruction tuning on ProLLaMA obtained from the previous section.
We train ProLLaMA on the instruction dataset mentioned in Section 3.1:
Here, $\Theta$ denotes the parameters to be optimized, $\mathcal{L}$ the loss function, and $\mathcal{D}$ the dataset. $\boldsymbol{u}$ denotes the instruction and the input of one instance. For brevity, when we refer to the instruction in the following text, it includes the input as well. $\boldsymbol{x}=\{x_{0},x_{1},\ldots,x_{n-1}\}$ denotes the output, where $x_{i}$ is the $i$-th token of the output.
Since causal language modeling is employed, we need to combine Equation 1 with Equation 3:
$\mathcal{L}(\Theta)=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D}}\left[-\sum_{i}\log p(x_{i}|\boldsymbol{u},x_{0},x_{1},\ldots,x_{i-1};\Theta)\right]$ (2)
Equation 2 is the optimization objective for instruction tuning of ProLLaMA. $\boldsymbol{u}$ is not involved in the loss calculation, whereas $\boldsymbol{x}$ is. This is because the latter is the output part, where ensuring its quality of generation is crucial. The former, $\boldsymbol{u}$, only needs to be understood by the model. In the instruction tuning stage, we exclusively train LoRA at a lower rank than specified in Section 3.3.
In summary, through instruction tuning, we have made ProLLaMA capable of following instructions and performing multiple tasks. Consequently, we have addressed the problems mentioned in Section 1: the lack of instruction following and the lack of multi-task capabilities.
In addition, the tasks that ProLLaMA can perform depend on the tasks included in the instruction dataset. Appendix A.7 shows that by training on an additional instruction dataset, ProLLaMA also performs well in predicting protein solubility.
Table 2: Comparison of proteins generated by different models. Our ProLLaMA achieves the best performance on pLDDT, TM-score, and RMSD metrics, and is second-best in SC-Perp, demonstrating ProLLaMA excels in de novo protein design. *: We list ProGen2 and ProGen as the same item, referring to Appendix A.6 for explanation. AE: Auto-Encoder. AR: Auto-Regressive.